# Macroeconomic Spillovers of Weather Shocks across U.S. States

### v2

@@ 4.2 The role of economic linkages and spatial adjacency
Figure 8 reports the IRFs of MECI calculated based on geographical proximity. Cross-region effects of weather shocks are almost entirely wiped out, and intra-regional effects are substantially reduced, relative to the baseline results in Figure 7.111111Figures 7 and 8 are reported in tabular form in the online Appendix.
@@ 4.3 The role of parameter heterogeneity and disaggregation
A spatial dynamic panel model. In the GVAR specification parameters are heterogeneous across states. To highlight the importance of such heterogeneity, we compare the baseline GVAR results with the estimates of a spatial dynamic model (i.e., a model with both space and time lags), in which all states share the same parameters:
$y_{it}=p\tilde{W_{i}}y_{it}+\gamma y_{i,t-1}+\rho\tilde{W_{i}}y_{i,t-1}+\beta s_{it}+c_{i}+e_{it}$ (8)
where $\tilde{W_{i}}$ is a $N$-dimensional vector of trade weights, $p,\gamma,\rho,\beta$ are parameters, $c_{i}$ denotes state fixed effects and $e_{it}$ is the error term. We use a tilde to distinguish the vector of weights on which we rely here from the $2\times N$ matrix used in (4), but the bilateral weights are exactly the same in the two models. Yu et al., (2008) have developed quasi-maximum likelihood estimators for spatial dynamic panel models of the form (8). In Panel (a) of Table 1, we report bias-corrected estimates of model (8), using 1000 iterations of the Yu et al., (2008) algorithm. The coefficient on weather shocks is around $-0.073$ and statistically distinguishable from zero.
@@
$\tilde{W_{i}}y_{it}$ 0.6910*** 0.0094
$y_{i,t-1}$ 0.1586*** 0.0074
$\tilde{W_{i}}y_{i,t-1}$0.0309*** 0.0119
$s_{it}$0.0734*** 0.0099
@@
$\theta_{i}$ $-0.066$ $0.11$ $-0.61$ $0.065$
Notes: Panel (a) reports estimates of a spatial dynamic panel model. The estimates of an autoregressive distributed lag (ARDL) model for the aggregate U.S. economy are shown in panel (b). Panel (c) displays the average point estimate of coefficient $\theta_{i}$ (i.e. the impact effect of weather shocks) across the state-specific ARX* models (3) composing the GVAR, along with the standard deviation of the estimates across states, and the minimum and the maximum estimates. In panels (a) and (b), ∗∗∗ indicates 1% significance, the other coefficients in panel (b) are not significant at the 10% level.
A country-wide regression for the U.S. Next, we consider a model that does not feature geographical disaggregation. Specifically, we estimate an autoregressive distributed lag model (ARDL) for the aggregate U.S. economy at the monthly frequency. As shown in Panel (b) of Table 1, the effect of weather shocks is negative but statistically indistinguishable from zero. We conclude that disaggregation matters, when it comes to estimating the economic impact of weather events.
Finally, for comparison, panel (c) of Table 1 reports the average point estimate of the impact effect of weather shocks, $\theta_{i}$, across the state-specific ARX* models (3) composing the GVAR, along with the standard deviation of the estimates across states, the minimum and the maximum estimates. The average estimate of $\theta_{i}$ in the GVAR model ($-0.066$) is similar to the coefficient on weather shocks in the spatial dynamic model ($-0.073$). However, the dynamic panel model fails to capture the heterogeneity of the response to weather shocks across states. As highlighted in panel (c), the estimated effects of weather shocks vary substantially across states in the GVAR model: the point estimates of coefficient $\theta_{i}$ have a standard deviation of $0.11$ across states, with values in the range from $-0.61$ to $0.065$ (for Louisiana and Maine, respectively).
4.4 The economic effects of specific weather-related disasters
The adverse weather conditions during winter and tropical storms - hurricanes in particular - are often the headline news due to their severe impacts on economic activity and supply chains. We thus focus on disasters triggered by such weather shocks in the climate regions where they happen more often.
In the case of shocks associated with adverse winter weather conditions, these are concentrated in the Northeast, Ohio Valley, and Southeast regions. As shown in the upper panel of Figure 9, shocks originating in any of these regions have significant negative impacts in all of the other regions. We can note that the Upper Midwest is particularly affected by shocks originating in the Northeast, Ohio Valley, and Southeast regions. This can be rationalized by looking at Figure 5, which shows that these regions are not only geographically close but also economically tightly linked through trade flows to the Upper Midwest.
As for shocks triggered by tropical storms, these are concentrated in the Northeast, Southeast, and South of the U.S. The lower panel of Figure 9 highlights that in this case only shocks originating in the South of the U.S. have negative and statistically significant impacts on the rest of the U.S. The magnitude of the response is larger for the neighboring regions and for the Northeast and the Upper Midwest than for western regions. This can be in part explained by the fact that the single most disruptive weather event in the recent history of the U.S. – Hurricane Katrina – has impacted the South of the country.
Figure 9: IRFs of Monthly Economic Conditions Indicators at 12-month horizon: Winter Weather and Tropical Storms
(a) Winter Weather
(b) Tropical Storms
Notes: see notes to Figure 7.
@@ 4.5 Indirect effects of weather shocks on local activity
To quantify indirect effects of weather shocks on local activity for each state, we compare the IRFs derived from the GVAR model with alternative IRFs derived using state-specific ARDL ($ARX^{\star}$) models and assuming that RoUS variables are not affected by changes in state-specific variables. The difference between the IRFs in the two cases captures “second-round effects”: a weather shock in state $i$ causes economic adjustment in the other states, which in turn feeds back into state $i$.
Figure 10 isolates second-round effects by comparing GVAR and ARDL estimates of the response of MECI to a weather shock after one year. In absolute terms, and independently of the model being used, these effects are especially strong in Louisiana (LA) and Illinois (IL). While the first response is chiefly driven by the impact of Hurricane Katrina in 2005, the second is associated with the Great Flood of 1993, which caused one of the most extensive financial losses and widespread devastation in modern U.S. history. More generally, we note that, as expected, GVAR estimates have wider confidence intervals but also tend to be greater in absolute value, hence reinforcing the conclusion that spillovers do matter in assessing the impacts of weather shocks. As a back-of-the-envelope calculation, the legend of the figure indicates that in absolute terms, both the average and median responses from the GVAR are larger than those from the ARDL model.
@@
Notes: For each state, the figure reports the bootstrap mean estimates of 12-month-ahead responses of state-level MECI to a simulated weather shock in the same state, provided by the U.S. GVAR model (blue squares) and by state-specific ARDL ($ARX^{\ast}$) models in which the rest of the U.S. economy is exogenous (red circles). The bars denote the 10% and 90% percentile from the bootstrap distribution.
5 Conclusions

### v3

@@ 4.2 The role of economic linkages and spatial proximity
Figure 7 reports the IRFs of MECI calculated by considering geographical proximity. Cross-region effects of weather shocks are almost entirely wiped out, and intra-regional effects are substantially reduced, relative to the baseline results in Figure 6.111111Figures 5-7 are reported in tabular form in the online Appendix.
@@ 4.3 The role of parameter heterogeneity and disaggregation
A spatial dynamic panel model. In the GVAR specification, the parameters are heterogeneous across states. To highlight the importance of such heterogeneity, we compare the baseline GVAR results with the estimates of a spatial dynamic model (i.e., a model with both space and time lags) in which all states share the same parameters:
$y_{it}=p\mathbf{\widetilde{W}}_{i}\mathbf{y}_{t}+\gamma y_{i,t-1}+\rho\mathbf{\widetilde{W}}_{i}\mathbf{y}_{t-1}+\beta s_{it}+c_{i}+e_{it}$ (8)
where $\mathbf{\widetilde{W}}_{i}$ denotes an $N$-dimensional row vector of trade weights, $p,\gamma,\rho,\beta$ are homogeneous scalar parameters, $c_{i}$ denotes state fixed effects and $e_{it}$ is an error term. We use a tilde to distinguish the weight vector employed here from the $2\times N$ weight matrix used in (4), but the bilateral weights are exactly the same in the two models (i.e., $y_{it}^{\ast}=\mathbf{\widetilde{W}}_{i}\mathbf{y}_{t}$). Yu et al., (2008) have developed quasi-maximum likelihood estimators for spatial dynamic panel models of the form (8). In panel (a) of Table 1, we report bias-corrected estimates of model (8), using the Yu et al., (2008) method. The coefficient on weather shocks is around 0.074-0.074 and statistically different from zero.
@@
$\mathbf{\widetilde{W}}_{i}\mathbf{y}_{t}$ 0.7109*** 0.0089
$y_{i,t-1}$ 0.1579*** 0.0074
$\mathbf{\widetilde{W}}_{i}\mathbf{y}_{t-1}$ 0.0392-0.0392*** 0.0121
$s_{it}$ 0.0737-0.0737*** 0.0099
@@
$\theta_{i}$ 0.067-0.067 $0.11$ 0.61-0.61 $0.065$
Notes: Panel (a) reports estimates of a spatial dynamic panel model. The estimates of an autoregressive distributed lag (ARDL) model for the aggregate U.S. economy are shown in panel (b). Panel (c) displays the average point estimate of coefficient $\theta_{i}$ (i.e., the impact effect of weather shocks) across the state-specific ARX* models (3) composing the GVAR, along with the standard deviation of the estimates across states, and the minimum and maximum estimates. In panels (a) and (b), ∗∗∗ indicates 1% significance, the other coefficients in panel (b) are not significant at the 10% level.
A country-wide regression for the U.S. Next, we consider a model that does not feature geographical disaggregation. Specifically, we estimate an autoregressive distributed lag (ARDL) model for the aggregate U.S. economy at the monthly frequency. As shown in panel (b) of Table 1, the effect of weather shocks is negative but statistically indistinguishable from zero. We conclude that disaggregation matters, when it comes to estimating the economic impact of weather events.
Finally, for comparison, panel (c) of Table 1 reports the average point estimate of the on-impact effect of weather shocks, $\theta_{i}$, across the state-specific ARX* models (3) composing the GVAR, along with the standard deviation of the estimates across states, as well as the minimum and maximum estimates. The average estimate of $\theta_{i}$ in the GVAR model (0.067-0.067) is similar to the coefficient on weather shocks in the spatial dynamic model (0.074-0.074). However, the dynamic panel model fails to capture the heterogeneity of the response to weather shocks across states. As highlighted in panel (c), the estimated effects of weather shocks vary substantially across states in the GVAR model: the point estimates of coefficient $\theta_{i}$ have a standard deviation of $0.11$ across states, with values in the range from 0.61-0.61 to $0.065$ (for Louisiana and Maine, respectively).
@@ 4.4 Indirect effects of weather shocks on local activity
To quantify indirect effects of weather shocks on local activity for each state, we compare the IRFs derived from the GVAR model with alternative IRFs derived using state-specific ARDL (ARX*) models and assuming that changes in state-specific variables do not affect other states. The difference between the IRFs in the two cases captures “second-round effects”: a weather shock in state $i$ causes economic adjustment in the other states, which in turn feeds back into state $i$.
Figure 8 isolates second-round effects by comparing GVAR and ARDL estimates of the response of MECI to a weather shock after one year. In absolute terms, these effects are especially strong in Louisiana (LA) and Illinois (IL). The results for the former are largely driven by the devastating impact of Hurricane Katrina in 2005, while those for the latter are especially associated with extreme winter weather events, such as the Blizzard of 1999.
More generally, we note that, as expected, GVAR estimates have wider confidence intervals but also tend to be greater in absolute value, hence reinforcing the conclusion that spillovers do matter in assessing the impact of weather shocks. As a back-of-the-envelope calculation, the legend of the figure indicates that, in absolute terms, both the average and median responses from the GVAR are larger than those from the ARDL models.
@@
Notes: For each state, the figure reports the bootstrap mean estimates of 12-month-ahead responses of state-level MECI to a simulated weather shock in the same state, provided by the U.S. GVAR model (blue squares) and by state-specific ARDL ($ARX^{\ast}$) models in which the rest of the U.S. economy is exogenous (red circles). The bars indicate the 10th and 90th percentiles from the bootstrap distribution.
5 Further results and robustness checks
In this section, we present a number of additional results and robustness checks. For the sake of brevity, the tables and figures underlying these results are presented in the online Appendix, to which the reader is referred.
5.1 Alternative proxies of weather related disasters
In this section, we consider two alternative proxies of weather shocks: one based on ex-ante impact estimates generated using the CLIMADA modelling framework, and the other on the Actuaries Climate Index, which provides a composite measure of climate extremes – including temperature anomalies, heavy precipitation, prolonged dry spells, high winds, and sea level changes – relevant for assessing economic vulnerability.
CLIMADA. The first alternative proxy is based on a measure of the ex-ante estimated impact of weather events, based on the physical features of the events and the economic exposure of the affected areas. To this end, we make use of CLIMADA (CLIMate ADAptation), a free and open-source software framework [^12] For more information on the current version, see the project website at https://climada.ethz.ch/., specifically designed to quantify the impacts of natural hazards and to assess climate risk (Aznar-Siguan and Bresch, , 2019). CLIMADA allows making projections of the damages from natural events at high geographical resolutions, by integrating data on hazards, data on exposures, and measures of vulnerability in the form of impact functions mapping the physical severity of hazards into economic losses. [^13] Specifically, an impact function determines the percentage of damage to an exposed asset as a function of hazard intensity. In the modelling community, this is often referred to as a vulnerability curve. Further details on the underlying data and CLIMADA modules are provided in the Appendix.
For each U.S. state, we use CLIMADA to generate a time series of projected losses caused by several types of weather events, closely matching those considered in the FEMA classification. Specifically, we calculate losses caused by tropical cyclones, other types of storms, river floods and wildfires. We first calculate the daily losses caused by each single event using granular geographic information (i.e., gridded data at a high resolution) on the physical characteristics of events and on the economic exposures. Then, we aggregate the estimated losses at the state level, at the monthly frequency, and across different events, to have a time series of losses that can be used in the GVAR in place of the FEMA-based proxy.
Figure 9(a) provides an example for a representative state, Florida. The leftmost figure plots the gridded economic exposures based on nightlights and population, the central figure plots the maximum intensity at each location of a major event, namely Hurricane Andrew of August 1992 (the intensity is measured by wind speed in m/s), and the figure on the right plots the time series of projected losses, reported as percentages of total state assets. Of course, the largest exposures are associated with main urban centers: Miami, Orlando, Tampa, Jacksonville. The peaks in the projected impact series are associated with major hurricanes, like Hurricane Andrew (August 1992) or Hurricane Wilma (October 2005).
Figure 9: Gridded impact analysis in CLIMADA: an example with Florida
(a) Notes: The figure on the left shows gridded economic exposure, based on night-time lighting and population density. Each dot represents a centroid, and the color scale indicates the level of exposure. For illustrative purposes, the absolute exposure scale is set using total household wealth, as provided by CLIMADA based on Credit Suisse data. The central figure shows the maximum intensity of Hurricane Andrew, measured by wind speed in meters per second (m/s), in Florida in August 1992, based on IBTrACS data. The figure on the right shows the CLIMADA-based indicator of weather shocks for Florida, in % of total state assets, from 1990m1 to 2019m12. This indicator is constructed using high-resolution data on tropical cyclones, other storms and river floods.
Using this approach, we find that weather shocks that generate significant effects are those hitting the South and Southeast regions, while shocks in other regions have generally non-significant impacts. While the results for the South and Southeast regions are totally in line with our baseline results based on the FEMA proxy, the main differences (in particular, the insignificant impact of shocks in the Northeast) appear at least in part due to the different way of dealing with winter weather events, which in CLIMADA we capture only to the extent that they are associated with storm activity.
Actuaries Climate Index (ACI). We compare our results also with those obtained using the Actuaries Climate Index (ACI) as an alternative proxy of weather related disasters (see, e.g., Kim et al., , 2025). The ACI measures the frequency of extreme weather events and the extent of sea level change on a monthly basis, condensing information related to six variables: high and low temperatures, heavy precipitation, consecutive dry days, high winds, and sea level.
Figure 10 plots both the ACI and our proxy for weather-related disasters for the U.S. Visual inspection of this plot reveals that the two time series exhibit very different patterns. In particular, the ACI shows a mean shift at the beginning of 1995; this is consistent with the empirical evidence in Kim et al., (2025), which highlights that both the mean and variability of the ACI are on average higher in recent years than in the past. Such behavior of the series suggests that the frequency of severe weather conditions, as captured by the ACI, has gradually increased over time. In contrast, our proxy for weather disasters does not capture the variation in the frequency of such weather conditions, but is more like a dummy for the occurrence of extreme events.
Figure 10: Weather Related Disasters and the Actuaries Climate Index (ACI)
Notes: The figure plots the (unsmoothed) ACI for the U.S., along with our FEMA-based indicator. Data on ACI are sourced from https://actuariesclimateindex.org
The ACI is not available at the state level, but for seven regions: Alaska (ALA), Central East Atlantic (CEA), Central West Pacific (CWP), Midwest (MID), Southeast Atlantic (SEA), Southern Plains (SPL), Southwest Pacific (SWP).
After properly aggregating state-level data at ACI regional level, we estimate a GVAR model using alternatively our measure of weather-related disasters and the ACI index as exogenous variables. In the case of our proxy, the estimated effects of weather shocks remain negative and generally significant at the 10% (one-sided), with the only exceptions of shocks hitting the SPL and SWP regions. When we use ACI, significance at the 10% is achieved only for shocks hitting the CWP region. Overall, the results suggest that focusing on emergency declarations triggered by weather-related disasters may provide sharper identification of disruptive weather shocks, compared to indicators of abnormal climate conditions.
5.2 On the transmission channels of weather shocks
In this section, we discuss the economic channels through which extreme weather events affect local economic activity and generate cross-border spillover effects.
MECI components. We start by re-estimating the GVAR model using, alternatively, the six sub-indices of the aggregate MECI indicators: (i) mobility, (ii) labor market, (iii) real activity, (iv) expectations, (v) households, (vi) financials. The complete set of results as well as details on the MECI sub-indices are reported in the Appendix.
The component of MECI that is associated with the strongest effects of weather shocks (both local and cross-border) and with high statistical significance in most regions is the “labor market” component. Another component that shows statistical significance in several regions is “mobility”, although in this case the magnitude of the impact and, more generally, the contribution of the component to total MECI are quite small. The responses of the “real activity”, “households”, and “expectations” components are non-negligible only in the case of shocks hitting the South, while the effects on the “financials” component are invariably non-significant.
Labor-market indicators. Next, to characterize in more detail the predominant labor-market channel, we study the responses of two specific indicators of labor activities: the number of hours worked in manufacturing and the growth rate of total employment. Both indicators generally drop after a weather-related disaster, both locally and in regions not directly hit. In the Appendix we also consider the impact of different types of weather events on employment. Winter weather, storms and floods have a significant negative impact in several regions, while tropical storms affect the Southeast in particular. Droughts and fires have no significant impact.
Weather shocks as supply-side shocks. Overall, the previous results, indicating a substantial impact of extreme weather on labor market conditions and mobility and, at the same time, a limited impact on typically demand-related components like “households” and “expectations”, suggest a central role of supply-side disruptions. To test that weather shocks are mainly supply-side shocks, we look at movements in consumer prices. Unfortunately, monthly inflation data are not available at the state level. However, the U.S. Bureau of Labor Statistics (BLS) publishes monthly data on 12-month inflation rates for four macro-regions from September 2004. The four regions are Northeast, Midwest, South, and West, and cover the entire U.S. territory.
A descriptive analysis of the data suggests that, following major weather-related disasters, inflation rates tend to increase across the entire U.S., regardless of where the event occurred. For instance, in August-September 2005, following Hurricane Katrina, the 12-month inflation rate increased by 2.4 percentage points in the Northeast, 2 in the Midwest, 1.8 in the South, and 1.1 in the West. Analogously, in April-May 2011, following the so-called Super Outbreak, the inflation rate increased by 0.7 percentage points in the Northeast, 1.1 in the Midwest, 1.2 in the South, and 0.6 in the West.
To test more formally that all regional inflation rates tend to increase in response to weather shocks, we estimate a VARX model for the four regional inflation rates, including our FEMA-based weather shock variable for the aggregate U.S. territory as a common exogenous variable. [^14] To control for the effects of aggregate economic activity and monetary policy on inflation, we also include the (12-month) growth rate of U.S.-wide industrial production and the Fed funds rate as two additional common exogenous variables. To account for potential price stickiness, we include both contemporaneous and lagged values of the exogenous variables. The lag order of the VARX is set to one, as suggested by the BIC. Thus, we consider the responses of regional prices to extreme weather events, regardless of where the events occurred.
The results confirm generalized price increases across regions, especially significant in the Midwest and in the South. This might suggest that the high degree of market integration across the U.S., as reflected in the strong co-movement of regional consumer price inflation rates, may act as a powerful mechanism underlying the cross-state transmission of shocks, thus helping explain the significant spillover effects found in our main results.
To sum up, our analysis so far suggests two main conclusions regarding the channels of transmission of weather-related disasters: (i) weather shocks mainly affect the local economy through supply and labor market disruptions; (ii) as a result of market integration, supply contractions appear to spread across the U.S., even if not all regions are directly hit by extreme events. However, further evidence will be needed to corroborate this interpretation of the results and to study supply-side effects. While we mostly leave this to future research, below we further explore supply-side contagion as a source of spillovers.
Inspecting the mechanisms behind spillovers. Lastly, since our main results reveal that geographic spillovers from severe weather events are strongly tied to economic interconnections between states, we now seek to further investigate the underlying causes and transmission mechanisms of these spillover effects.
The literature highlights the critical role of supply-chain (input-output) relationships in the transmission of natural disaster shocks. Barrot and Sauvagnat, (2016) show that disasters disrupt production networks in the U.S. by affecting supplier firms, thereby causing significant indirect losses to downstream customer firms – even in unaffected areas – due to costly input substitution. Similarly, Hallegatte, (2008, 2014) provides evidence of substantial national economic losses from disasters like Hurricane Katrina via sectoral linkages. International spillovers are also observed, as shown by Feng et al., (2025), Forslid and Sanctuary, (2023), Kashiwagi et al., (2021), and Boehm et al., (2019), who find that disasters impact countries indirectly through global supply chains. Building on this literature, we present two quantitative assessments – each employing a distinct methodological approach – of the origins and magnitude of spillover effects. Methodological details and full results are provided in the Appendix.
In the first exercise, we return to our GVAR model to investigate whether economic spillovers are more strongly linked to trade in intermediate goods rather than final goods, as suggested by earlier studies. Although the available data do not explicitly classify interstate shipments by final or intermediate use, we rely on the sectoral information provided by the U.S. Commodity Flow Survey to make this distinction. By combining these data with input-output tables from the U.S. Bureau of Economic Analysis, we classify sectors based on whether the majority of their output is used as intermediate inputs or final goods.
Using this classification, we build two separate trade-weight matrices, $\mathbf{W}$: one reflecting trade in intermediate-good-intensive sectors and the other for final-good-intensive sectors. We then re-estimate the GVAR model using either matrix separately. The results indicate that spillovers are stronger when focusing on trade in intermediate goods, thus reinforcing the idea that input linkages across regions and sectors play a key role in amplifying the effects of economic shocks.
The second exercise explores the indirect economic consequences of natural disasters by focusing on how disruptions in one part of the economy can ripple across sectors. Using the CLIMADA software, the analysis focuses on tropical cyclones in the United States from 1990 to 2019. We estimate both the immediate (direct) damages in affected areas and the broader (indirect) nationwide impacts which arise from input-output relationships between economic sectors.
As shown in Figure A13 in the Appendix, indirect losses are substantial, nearly half the size of direct damages. For instance, in 2005, the year of Hurricane Katrina, direct losses from tropical cyclones are estimated at around 0.75% of GDP, while indirect losses add another 0.3%. Although this is a static analysis that does not account for longer-term economic adjustments, it still underscores the significant role of supply-chain linkages in amplifying the economic costs of natural disasters.
5.3 The possible endogeneity of the weather shock indicator
The political influences on emergency declarations and the quantification of relief funding have been addressed in different studies, e.g., Garrett and Sobel, (2003). Our indicator of weather shocks, being based on emergency declarations, might contain information not strictly related to natural disasters, with possible concerns about endogeneity. [^15] We thank an anonymous referee for raising this interesting issue. Felbermayr and Gröschl, (2014) propose a dataset (denoted GeoMet) containing measures of the physical strength for a set of natural disasters. Such indicators, not based on monetary damages, solve the potential endogeneity problem, as well as, in studies relating to very different countries, the issue of higher damages being generally associated with richer countries. On the other hand, these indicators, based on different measurement scales, might give rise to aggregation issues, if the aim is to define a unique indicator of natural disasters. Moreover, particular types of natural disasters, like wildfires, are very difficult to measure with a single indicator.
To deal with this potential identification issue, we propose a robustness check based on a two-stage approach, where in the first stage our indicator of natural disasters is regressed, state by state, on the number of deaths and a dummy variable denoting whether the disaster is included in the U.S. Billion-Dollar Weather & Climate Disaster database. [^16] As the FEMA and the Billion-Dollar Weather & Climate Disaster databases are not perfectly synchronized, especially for events that occurred over long time spans (e.g., droughts, wildfires), or occurred across two distinct months, we preliminarily performed a deep data cleaning to match events in the two databases. All modifications, however, have been implemented when an appropriate documentation was found.
The results of the first stage regressions confirm that this exogenous external information significantly helps to capture the main features of the weather shock indicator. Even more important, in the second stage, where the weather shock indicator is replaced by the endogenous-free fitted values of the first stage, almost all the results described in Section 4.1 are confirmed, both in terms of the impact on the local economic activity measure and of the spillover effects.
To further address potential endogeneity, we also introduce an alternative instrumental variable based on extreme weather events. Our earlier approach relied on the number of deaths caused by such events, using it as a proxy for severity, while deliberately avoiding monetary measures of impact. We now refine this strategy by considering an instrument that is entirely independent of any type of impact – monetary or non-monetary, observed or expected. Rather than reflecting consequences, this instrument is grounded solely in the physical attributes of natural events and the ex-ante exposure of the affected areas.
To construct this alternative instrumental variable, we follow Bilal and Rossi-Hansberg, (2023) and build a storm indicator that reflects the physical severity of weather phenomena (specifically, maximum precipitation and wind speed—across U.S. states). The construction involves removing state and seasonal fixed effects from the raw weather variables, and isolating extreme realizations via residuals exceeding the 95th percentile. The final indicator, $D^{storm}_{it}$, captures whether a state experiences either unusually high precipitation or wind speed in a given month. To incorporate the economic relevance of these events, the storm variable is weighted using gridded economic activity data. [^17] The underlying climate data come from the Weighted Climate Dataset (WCD) developed by Gortan et al., (2024), which integrates socio-economic and climate metrics. Further details are provided in the Appendix. This storm indicator is then used as an instrument for our FEMA-based disaster variable in a two-stage least squares estimation within the GVAR framework. The results largely confirm our baseline findings, with significantly negative effects in the Southeast, South, and Ohio Valley regions, although effects in the Northeast are no longer statistically significant in this case.
5.4 Robustness checks
Nonlinearities. We rely on two different approaches to assess potential nonlinearities in the relationships between weather shocks and economic activity. First, we re-estimate the U.S. GVAR adding the square of the weather shock variable as an additional regressor. Second, we allow for the possibility that the negative impact of adverse weather events on a state’s economy may be amplified if the state has already experienced a weather shock in the recent past. To this end, we include an interaction term between our weather shock variable $s_{it}$ and a dummy variable indicating whether the state experienced a severe weather shock in the last year. In both cases, results remain qualitatively similar to those in Section 4.1.
Including state-level control variables. The inclusion of RoUS averages in state-specific ARX* models allows us to account for common factors affecting the aggregate U.S. economy. We now check whether our results are robust to the inclusion of state-specific variables that may affect state-level economic activity. We thus re-estimate the GVAR model including two additional regressors in each ARX* model: the government component of Gross State Product and state exports to the rest of the world. As in our baseline results, the effects of weather shocks are negative and at least one-sided 10% significant for shocks hitting the Northeast, Southeast, South and Northern Rockies and Plains regions. In addition, negative and significant results are now obtained also for the West and Northwest regions.
Alternative county weights. Our baseline weighting of weather events, based on the percentage of counties hit in a state, is intended to capture the intensity of disasters rather than the importance of individual counties. As a robustness check, however, we consider an alternative weighting scheme reflecting the different population levels of affected counties. Specifically, for each state, we scale the weather event dummy using the population of counties hit as a share of the state’s total population. Once again, full results are reported in the Appendix. The effects of weather shocks remain significantly negative for events hitting the Northeast and South regions, and close to significance for events hitting the Southeast. Thus, the results are broadly in line but weaker than the baseline. This appears consistent with the fact that productive establishments (such as factories), whose activities are disrupted by extreme weather events, are often located in areas that are not densely populated.
Further robustness checks. We report additional robustness checks in the Appendix. These include re-estimating the model over a shorter sample that ends before the Global Financial Crisis of 2007–2009, assessing the potential endogeneity of the trade weights, and examining their time constancy. We also test the robustness of our results to alternative specifications of the trade-weight matrices.
6 Conclusions and policy implications